Design a circuit find the sum of 2 input voltages V1 and V2 op-amp may be used in inverting mode


1
Expert's answer
2021-06-11T06:11:15-0400

Employing the circuit below to get "V_{out}=-K_1*V_1+K_2*V_2" for a reasonable gains with positive values of K1 and K2 .




"V_{out}=-V_1(\frac{R_3}{R_1})+V_2(\frac{R_4}{R_2+R_4})(\frac{R_1+R_3}{R_1})"

Assuming K1=1 and K2=2 are your requirements, define R1=R3 with a suitable value.


With this limitation, you must construct K2 with R2=0, thus link V2 straight to the (+) input, rendering R4 obsolete and allowing R4 to be deleted.


In a nutshell, the answer is to connect V2 to the (+) input terminal directly and set R1=R3.


In reality, this circuit is frequently used with R1=R2 and R3=R4 to provide a balanced subtractor with gain R3/R1=R4/R3 in each input, i.e., Vout=(R3/R1)* (V2-V1).